Two crossed belts on pulleys of diameters 3.6 m and 2.4 m connect two parallel shafts with centres 4.2 meters apart. The maximum tension in the belts is limited to 1200 N and friction between the belts and the pulley, μ = 0.26. The smaller pulley has a speed of 300 rev/min.

2.1. Find the power that can be transmitted. (8)

2.2. What would be transmitted if open belts were used. (8)


1
Expert's answer
2021-07-13T10:42:08-0400

1)

"P_1=\mu F l\nu\frac {d_2}{d_1}=9.8~kW,"

2)

"P_2=P_1\frac{d_1^2}{d_2^2}=4.4~kW."